For each of the following production functions, determine whether returns to scale are decreasing, constant, or increasing,


a. Q = 2K + 3L + KL

b. Q = 20K0.6 L0.5

c. Q = 100 + 3K + 2L

d. Q = 5K0.459 L0.541

e. Q = 5K0.37 L0.56

f. Q = K/L


1
Expert's answer
2022-03-29T12:13:03-0400

Basically, the returns to scale refers to how much output changes given a proportional change in all inputs, where all the inputs change by the same factor.

To obtain the next output, we shall multiply all our functions by a factor z


a. Q = 2K + 3L + KL


"Let \ud835\udc44_{0}=\ud835\udc39(\ud835\udc3e,\ud835\udc3f)=2\ud835\udc3e+3\ud835\udc3f+\ud835\udc3e\ud835\udc3f"


multiply it by factor 𝑧 and call it 𝑄1

"\ud835\udc441=\ud835\udc39(\ud835\udc67\ud835\udc3e,\ud835\udc67\ud835\udc3f)=2(\ud835\udc67\ud835\udc3e)+3(\ud835\udc67\ud835\udc3f)+(\ud835\udc67\ud835\udc3e)(\ud835\udc67\ud835\udc3f)=\ud835\udc67(2\ud835\udc3e+3\ud835\udc3f+\ud835\udc67\ud835\udc3e\ud835\udc3f)"

"\ud835\udc67(2\ud835\udc3e+3\ud835\udc3f+\ud835\udc67\ud835\udc3e\ud835\udc3f)>\ud835\udc67(2\ud835\udc3e+3\ud835\udc3f+\ud835\udc3e\ud835\udc3f)"

"2\ud835\udc3e+3\ud835\udc3f+\ud835\udc67\ud835\udc3e\ud835\udc3f>2\ud835\udc3e+3\ud835\udc3f+\ud835\udc3e\ud835\udc3f"

"\ud835\udc67\ud835\udc3e\ud835\udc3f>\ud835\udc3e\ud835\udc3f"

"\ud835\udc67>1."

This indicates increasing returns to scale. 


b. Q = 20K0.6 L0.5

"\ud835\udc441=\ud835\udc39(\ud835\udc67\ud835\udc3e,\ud835\udc67\ud835\udc3f)=20(\ud835\udc67\ud835\udc3e)^{0.6}(\ud835\udc67\ud835\udc3f)^{0.5}=\ud835\udc67^{0.6}\ud835\udc67^{0.5}20\ud835\udc3e^{0.6}\ud835\udc3f^{0.5}=\ud835\udc67^{1.1}\ud835\udc44_{0 }"

"\ud835\udc39(\ud835\udc67\ud835\udc3e,\ud835\udc67\ud835\udc3f)> \ud835\udc67\ud835\udc39(\ud835\udc3e,\ud835\udc3f)."

This function expresses increasing returns to scale


c. Q = 100 + 3K + 2L

"\ud835\udc441=\ud835\udc39(\ud835\udc67\ud835\udc3e,\ud835\udc67\ud835\udc3f)=100+3(\ud835\udc67\ud835\udc3e)+2(\ud835\udc67\ud835\udc3f)"

In this case;

"\ud835\udc39(\ud835\udc67\ud835\udc3e,\ud835\udc67\ud835\udc3f)<\ud835\udc67\ud835\udc39(\ud835\udc3e,\ud835\udc3f)"

"100+3(\ud835\udc67\ud835\udc3e)+2(\ud835\udc67\ud835\udc3f)<100\ud835\udc67+3(\ud835\udc67\ud835\udc3e)+2(\ud835\udc67\ud835\udc3f)"  This production function represents decreasing returns to scale.


d. Q = 5K0.459 L0.541

Let "\ud835\udc44_{0}=\ud835\udc39(\ud835\udc3e,\ud835\udc3f)=5\ud835\udc3e^{\ud835\udc4e}\ud835\udc3f^{\ud835\udc4f}" , where 𝑎+𝑏=1, be the initial production function, then after multiplying it by factor 𝑧 we obtain: 

"\ud835\udc44_{1}=\ud835\udc39(\ud835\udc67\ud835\udc3e,\ud835\udc67\ud835\udc3f)=5(\ud835\udc67\ud835\udc3e)^\ud835\udc4e(\ud835\udc67\ud835\udc3f)^\ud835\udc4f=\ud835\udc67^\ud835\udc4e\ud835\udc67^\ud835\udc4f5\ud835\udc3e^\ud835\udc4e\ud835\udc3f^\ud835\udc4f=\ud835\udc67^{\ud835\udc4e+\ud835\udc4f}\ud835\udc44_{0}=\ud835\udc67\ud835\udc44_{0}"

"\ud835\udc39(\ud835\udc67\ud835\udc3e,\ud835\udc67\ud835\udc3f)=\ud835\udc67\ud835\udc39(\ud835\udc3e,\ud835\udc3f)."